A long copper wire with a radius of 2.0 mm carries a uniform linear charge density of λ = 6.2x10-8 C/m. Find the magnitude of the electric field at a distance of 10 cm from the wire.

1
Expert's answer
2022-03-09T10:32:42-0500
"E=\frac{\lambda}{2\pi\epsilon_0 r}"

"E=\frac{6.2*10^{-8}\rm\: C/m}{2\pi*8.85*10^{-12}\:\rm F/m*0.102\: m}=1.1*10^4\:\rm N/C"